In an arithmetic series, the terms of the series are equally spread out. For example, in
1 + 5 + 9 + 13 + 17, consecutive terms are 4 apart. If the first term in an arithmetic series is
3, the last term is 136, and the sum is 1,390, what are the first 3 terms?

Answer by edjones(8007)

n/2(a[1]+a[n])=S[n]
n/2 * (3+136)=1390
n/2 * 139 = 1390
278n=2780 Multiply each side by 2.
n=10
.
136-3=133
133/9 common difference=d
.
c=a[1]-d
c= 3 - 133/9 = -106/9
a[n]=dn+c
a[2]= 133/9 * 2 -106/9 = 160/9
.
a[3]= 133/9 * 3 -106/9 = 293/9
.
Ed